OpenStudy (anonymous):

solve by factoring: -x^4+2x^3+8x^2=0

OpenStudy (anonymous):

first factor out -1 to get -x`2(x~2n +x +8)

Directrix (directrix):

-x^4+2x^3+8x^2=0 -x^2( x^2 - 2x + 8) = 0 Now, factor ( x^2 - 2x + 8).